How many gallons of a 15% salt solution need to be mixed with a 35% salt solution to obtain 8 gallons of 30% salt solution?
Write a system of linear equation that represents this problem situation. Do not solve the system.

Answer by stanbon(75887)

How many gallons of a 15% salt solution need to be mixed with a 35% salt solution to obtain 8 gallons of 30% salt solution?
Write a system of linear equation that represents this problem situation
=============
Quantity Equation: x + y = 8
Salt Equation: 0.15x+0.35y = 0.30*8
========================================
Cheers,
Stan H.